elimination
5x + 2y = 26
7x + 6y = 14

Answer by mananth(16946)

5 x + 2 y = 26 .............1

7 x + 6 y = 14 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 1
-15 x -6 y = -78
7 x + 6 y = 14
Add the two equations
-8 x = -64
/ -8
x = 8
plug value of x in (1)
5 x + 2 y = 26
40 + 2 y = 26
2 y = 26 -40
2 y = -14
y = -7